OpenStudy (anonymous):

Find the y-intercept for the line whose equation is given below. -4x + 2y = 10

OpenStudy (hoblos):

(0,5)

OpenStudy (anonymous):

fill in 0 for x: -4(0)+2y=10 0+2y=10 y=10/2 y=5 (0,5)